Let f element R(alpha) on [a,b] where alpha is of bounded variation on [a,b] and let v(x) denote the total variation of f on [a,x] for each x in [a,b] and let v(a) =0 , show that | integral a to b f d alpha |less than or equal to integral a to b |f| dv less than or equal to M.v(b) where M is an upper bound for |f| on [a,b]
1
Expert's answer
2020-05-11T17:46:53-0400

Now we have "f\in\mathcal R(V)" on "[a,b]", hence "|f|\in\mathcal R(V)"

Also "|\Delta\alpha_i|\leqslant \Delta V_i" and we obtain

"\bigg| \sum\limits_i f(t_i) \Delta \alpha_i \bigg|\leqslant \sum\limits_i |f(t_i)||\Delta \alpha_i|\leqslant \sum\limits_i |f(t_i)|\Delta V_i"

and for the limits we have

"\bigg| \int\limits_a^b f \,d\alpha \bigg|\leqslant \int\limits_a^b|f|\,dV"

Next, "V" increases, so for any upper Stieltjes sum "US(|f|,V)" we have "US\leqslant M|V(b)-V(a)|=MV(b)." Hence, we have the same for the infinum of sums, i.e. for "\int\limits_a^b|f|\,dV" .

So, we showed that

"\bigg| \int\limits_a^b f \,d\alpha \bigg|\leqslant \int\limits_a^b|f|\,dV\leqslant MV(b)"

where "M=\sup\limits_{[a,b]} |f|"